OpenStudy (anonymous):

Solve the system of equations using the elimination method. Enter the answer as an ordered pair.-3y=x-5 x+5y=7

OpenStudy (anonymous):

if the number has a letter next to it means multiply

OpenStudy (anonymous):

so (9,25)?

OpenStudy (anonymous):

sorry but thats all can say.

OpenStudy (anonymous):

Let's do this...take: -3y=x-5 And subtract x from both sides. What does that give you?

OpenStudy (anonymous):

-3y=-5

OpenStudy (anonymous):

BOTH sides. You have to subtract x from the left side, too, not just the right side.

OpenStudy (anonymous):

5y=7

OpenStudy (anonymous):

-3y-5 and 5y=7

OpenStudy (anonymous):

No...we're just working with the first equation for now -3y=x-5 T subtract x from BOTH SIDES OF THE "=" SIGN, -3y-x=x-x-5 Which leaves -3y-x=-5 THAT is what is meant by subtracting x from both sides. Essentially, you just flip the sign and move it to the other side. Do you see that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay. Here is the step where we do "elimination". Take the equation you just worked on and add it to the second equation this way: 5y+x=7 -3y-x=-5 BTW, adding negative numbers is the same as subtracting them. What do you get when you do this?

OpenStudy (anonymous):

5y + -3y is the same as 5y - 3y, so you have 3 terms to solve: 5y - 3y = ? x - x = ? and on the other side, 7 - 5 = ?

OpenStudy (anonymous):

(2,2)?

OpenStudy (anonymous):

y=2 x=1 7-5=2

OpenStudy (anonymous):

7-5 does equal 2 5y-3y does not equal 2. It equals 2y. There's a MAJOR difference between "2" and "2y" x-x does not equal 1 What does x-x equal?

OpenStudy (anonymous):

5y-3y=2y x-x=0 7-5=2 SO: 5y+x=7 -3y-x=-5 ---------- 2y+0=2 Or simply: 2y=2 Now...can you tell me the value of y if 2y=2?